# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v3

@@ Appendix B Detailed Convergence Analysis
$\mathbb{E}\left[\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\right]=\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]+2\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\beta_{u}\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\nabla\mathcal{L}(\theta_{t})\|^{2}\right]$
$\leq$ $2\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\beta_{u}\sigma^{2},$
@@
$\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]\leq 4\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+4\beta_{u}\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}.$
@@
[30] Let $A$ and $B$ be two symmetric matrices, and $u$ obeys the Gaussian distribution, that is, $u\sim N(0,I_{d})$. Define $z=u^{\top}Au\cdot u^{\top}Bu$. The expectation of $z$ is:
@@ Appendix C Test Functions of the optimization trajectories
  • Function (a) [^1] Function (a) is from [26].: $f(x,y)=8(x-1)^{2}(1.3x^{2}+2x+1)+0.5(y-4)^{2}$
@@
Figure 9: Influence of estimation times per step. (left) 2D trajectories of gradient descent; (right) Corresponding loss curves.
@@
2:for $t=1,...,T$ do
3:     Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
4:     $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
6:     $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
8:     $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
10:     $\hat{\Sigma}^{-1}_{t-1}=(R_{t-1}*C_{t-1})/(1_{n}^{\top}*R_{t-1})$
16:     for $\theta_{i}\in\theta$ do
17:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
19:     end for
20:end for
21:
22:function PerturbParameter($\theta$, $\mu$, $R_{t}$, $C_{t}$, $s$)
24:     for $\theta_{i}\in\theta$ do
25:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
26:         $\hat{\Sigma}^{-1}_{t}=(R_{t}*C_{t})/(1_{n}^{\top}*R_{t})$
28:     end for
29:     return $\theta$
30:end function
Algorithm 3 HiZOO-L
Appendix D HiZOO Variants
D.1 HiZOO-multi
D.2 HiZOO-L
Due to the storage of Hessian, HiZOO introduces extra memory cost, which is equal to the size of the model parameters. To address this limitations, we propose HiZOO-L, the low-rank implementation for the storage of Hessian, motivated by Adafactor [41]. Details can be see in Algorithm 2. We conduct lots of experiments to show the performance of HiZOO-L, as results shown in Figure 10.
Figure 10: Accuracy comparison on OPT-13B (left) and Llama3 (right) between HiZOO and HiZOO-L.
@@ E.1 Detailed Experiments on RoBERTa-large
Here we plot more loss curves to compare with MeZO. As shown in Figure 11, we can see that HiZOO can greatly accelerate the training process over MeZO, which verifies the robustness of HiZOO.
E.2 Detailed results on OPT
We use the hyperparameters in Table 7 for HiZOO experiments on OPT. We also provide the relative loss curves of fine-tuning OPT family in Figure 12.
@@ E.2 Detailed results on OPT
E.3 Detailed results on Phi-2 and Llama3
We provide several loss curves of fine-tuning Phi-2(2.7B) and Llama3(8B) in Figure 13 and Figure 14.
@@ Appendix F Details about Memory Usage
Table 8: Memory usage on the MultiRC (average tokens=400) dataset. Results of ICL and full-parameter tuning are from MeZO[32].
@@ Appendix G Details about Wallclock Time Efficiency
Table 9: Wallclock time per step between MeZO and HiZOO. The increase in wallclock time per step for HiZOO compared to MeZO is less than 1.5 times across different model sizes. To avoid introducing additional overheads such as inter-GPU communication, results are measured on the same dataset (SST-2) and GPUs (80GB A100), with each result averaged over 100 steps. "BS" refers to batch size. For the relatively smaller RoBERTa-large model, we used a BS=64, while for models larger than 1B parameters, we used a BS=16.
@@ Appendix H Details about Ablation Experiments
H.1 Influence of Smooth Scale $\alpha_{t}$
We conducted experiments on SST-2, SST-5, MNLI datasets when fine-tuning RoBERTa-large to research the influence of smooth scale $\alpha_{t}$. Figure 15 shows that the value of $\alpha_{t}$ mainly affects the convergence speed of the model. Additionally, the best value of $\alpha_{t}$ will vary between different datasets.
@@ H.1 Influence of Smooth Scale $\alpha_{t}$
H.2 Experiments about Omitting $[-\Sigma^{-1}]$ term in Eq. (5)
@@
Figure 16: Experiment about the error generate by omitting the $[-\Sigma^{-1}]$ term in Eq (5). ’with’ means holding the term and ’without’ means omitting the term.

### v4

@@ Appendix B Detailed Convergence Analysis
$\mathbb{E}\left[\nabla^{\top}{\bm{L}}(\theta_{t})\Sigma_{t}\nabla{\bm{L}}(\theta_{t})\right]=\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]+2\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{t})-\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\beta_{u}\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{t})-\nabla{\bm{L}}(\theta_{t})\|^{2}\right]$
$\leq$ $2\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\beta_{u}\sigma^{2},$
@@
$\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]\leq 4\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+4\beta_{u}\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}.$
@@
(Magnus et al., 1978) Let $A$ and $B$ be two symmetric matrices, and $u$ obeys the Gaussian distribution, that is, $u\sim N(0,I_{d})$. Define $z=u^{\top}Au\cdot u^{\top}Bu$. The expectation of $z$ is:
@@ Appendix C Test Functions of the optimization trajectories
  • Function (a)Function (a) is from (Liu & Li, 2023).: $f(x,y)=8(x-1)^{2}(1.3x^{2}+2x+1)+0.5(y-4)^{2}$
@@
Appendix D HiZOO Variants
D.1 HiZOO-L
Due to the storage of Hessian, HiZOO introduces extra memory cost, which is equal to the size of the model parameters. To address this limitations, we propose HiZOO-L, the low-rank implementation for the storage of Hessian, motivated by Adafactor (Shazeer & Stern, 2018). Details can be see in Algorithm 30. We also visualize the loss curves of HiZOO and HiZOO-L in Figure 9 and find that on most datasets two algorithms perform closely. This also indicates that the estimation of Hessian in HiZOO may be sparse, so we encourage researchers to try other memory efficient algorithms to compress the Hessian.
2:for $t=1,...,T$ do
3:     Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
4:     $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
6:     $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
8:     $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
10:     $\hat{\Sigma}^{-1}_{t-1}=(R_{t-1}*C_{t-1})/(1_{n}^{\top}*R_{t-1})$
16:     for $\theta_{i}\in\theta$ do
17:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
19:     end for
20:end for
21:
22:function PerturbParameter($\theta$, $\mu$, $R_{t}$, $C_{t}$, $s$)
24:     for $\theta_{i}\in\theta$ do
25:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
26:         $\hat{\Sigma}^{-1}_{t}=(R_{t}*C_{t})/(1_{n}^{\top}*R_{t})$
28:     end for
29:     return $\theta$
30:end function
Algorithm 2 HiZOO-L
D.2 HiZOO-multi
Figure 9: Loss curves on Llama3 between HiZOO and HiZOO-L.
@@ E.1 Detailed Experiments on RoBERTa-large
Table 7: Experiments on RoBERTa-large (350M parameters, k=512). For MeZO we report the results we reproduced.
Task Type SST-2 SST-5 SNLI MNLI RTE TREC Average
—— sentiment —— —— natural language inference —— — topic —
Zero-shot 79.0 35.5 50.2 48.8 51.4 32.0 49.5
LP 91.3 ($\pm$0.5) 51.7 ($\pm$0.5) 80.9 ($\pm$1.0) 71.5 ($\pm$1.1) 73.1 ($\pm$1.5) 89.4 ($\pm$0.5) 76.3
FT 91.9 ($\pm$1.8) 47.5 ($\pm$1.9) 77.5 ($\pm$2.6) 70.0 ($\pm$2.3) 66.4 ($\pm$7.2) 85.0 ($\pm$2.5) 73.1
FT (LoRA) 91.4 ($\pm$1.7) 46.7 ($\pm$1.1) 74.9 ($\pm$4.3) 67.7 ($\pm$1.4) 66.1 ($\pm$3.5) 82.7 ($\pm$4.1) 71.6
FT (prefix) 91.9 ($\pm$1.0) 47.7 ($\pm$1.1) 77.2 ($\pm$1.3) 66.5 ($\pm$2.5) 66.6 ($\pm$2.0) 85.7 ($\pm$1.3) 72.6
MeZO 93.3 ($\pm$0.7) 53.2 ($\pm$1.4) 83.0 ($\pm$1.0) 78.3 ($\pm$0.5) 78.6 ($\pm$2.0) 94.3 ($\pm$1.3) 80.1
MeZO (LoRA) 90.5 ($\pm$0.6) 45.4($\pm$1.1) 64.6($\pm$1.2) 62.1($\pm$0.9) 61.1($\pm$1.8) 80.8($\pm$1.5) 67.4
MeZO (prefix) 93.3 ($\pm$0.1) 53.6 ($\pm$0.5) 82.9 ($\pm$1.1) 75.6 ($\pm$1.2) 77.2 ($\pm$0.8) 88.2 ($\pm$0.7) 78.4
HiZOO 95.5 ($\pm$0.4) 53.2 ($\pm$0.9) 82.6 ($\pm$0.7) 77.7 ($\pm$0.6) 80.0 ($\pm$ 1.5) 94.6 ($\pm$1.1) 80.6
HiZOO (LoRA) 91.7 ($\pm$0.3) 45.3 ($\pm$0.7) 76.5 ($\pm$0.3) 63.1 ($\pm$0.6) 70.4 ($\pm$1.4) 85.6 ($\pm$1.5) 72.1
HiZOO (prefix) 96.1 ($\pm$0.2) 54.2 ($\pm$0.4) 85.7 ($\pm$0.7) 79.7 ($\pm$1.0) 77.3 ($\pm$0.2) 93.9 ($\pm$0.6) 81.2
@@
Figure 11: Accuracy curves on RoBERTa-large between MeZO and HiZOO.
In Table 7 we show the full experiment results. Additionally, we plot more loss curves to compare with MeZO. As shown in Figure 10, we can see that HiZOO can greatly accelerate the training process over MeZO, which verifies the robustness of HiZOO.
E.2 Detailed results on various LLMs
We use the hyperparameters in Table 8 for HiZOO experiments on OPT. Full results for OPT-30B and OPT-66B are in Table 9. We also provide the relative loss curves of fine-tuning OPT family in Figure 14. We provide several loss curves of fine-tuning Phi-2(2.7B) and Llama3(8B) in Figure 14 and Figure 14.
@@ E.2 Detailed results on various LLMs
Table 9: Experiments on OPT-30B and OPT-66B(with $1000$ examples). The best results are highlighted in bold for better comparison. We highlight the best results between HiZOO and MeZO in bold to facilitate comparison.
Task SST-2 RTE WSC WIC
30B zero-shot 56.7 52.0 38.5 50.2
30B ICL 81.9 66.8 56.7 51.3
30B MeZO 90.6 66.4 63.5 48.9
30B MeZO(prefix) 87.5 66.1 55.8 59.1
30B HiZOO 90.3 69.3 63.5 53.4
30B HiZOO(prefix) 91.2 68.6 57.7 60.2
66B zero-shot 57.5 67.2 43.3 50.6
66B ICL 89.3 65.3 52.9 54.9
66B MeZO(prefix) 93.6 66.4 57.7 58.6
66B HiZOO(prefix) 93.6 71.5 60.6 61.1
@@ Appendix F Details about Memory Usage
Table 10: Memory usage on the MultiRC (average tokens=400) dataset. Results of ICL and full-parameter tuning are from MeZO(Malladi et al., 2023).
@@ Appendix G Details about Wallclock Time Efficiency
Table 11: Wallclock time per step between MeZO, HiZOO and HiZOO-L. The increase in wallclock time per step for HiZOO compared to MeZO is less than 1.5 times across different model sizes. To avoid introducing additional overheads such as inter-GPU communication, results are measured on the same dataset (SST-2) and GPUs (80GB A100), with each result averaged over 100 steps. "BS" refers to batch size. For the relatively smaller RoBERTa-large model, we used a BS=64, while for models larger than 1B parameters, we used a BS=16.
@@ Appendix H Details about Ablation Experiments
H.1 Influence of Smooth Scale $\alpha_{t}$ and number of estimation $n$ per step
We conducted experiments on SST-2, SST-5, MNLI datasets when fine-tuning RoBERTa-large to research the influence of smooth scale $\alpha_{t}$. Figure 15 shows that the value of $\alpha_{t}$ mainly affects the convergence speed of the model. Additionally, the best value of $\alpha_{t}$ will vary between different datasets. Figure 16 shows that the influence of the number of estimation $n$ per steps.
@@ H.1 Influence of Smooth Scale $\alpha_{t}$ and number of estimation $n$ per step
Figure 16: Influence of number of estimation per step. (left) 2D trajectories of gradient descent; (right) Corresponding loss curves.
H.2 Experiments about Omitting $[-\Sigma^{-1}]$ term in Eq. equation 4
@@
Figure 17: Experiment about the error generate by omitting the $[-\Sigma^{-1}]$ term in Eq equation 4. ’with’ means holding the term and ’without’ means omitting the term.